The age of Juan is 3/4 of the age of Pedro. After 6 years, Juan's age will be 4/5 of Pedro's age. How old are each of them now?
Answer by VFBundy(438)

Pedro's age = p
Juan's age = 3p/4

3p/4 + 6 = 4(p + 6)/5

3p/4 + 6 = (4p + 24)/5

5(3p/4 + 6) = 4p + 24

15p/4 + 30 = 4p + 24

4(15p/4 + 30) = 4(4p + 24)

15p + 120 = 16p + 96

-p = -24

p = 24

Pedro's age = p = 24
Juan's age = 3p/4 = 3(24)/4 = 72/4 = 18